OpenStudy (anonymous):

. How many cubic centimeters can a cigar box hold if its dimensions are 4 centimeters by 14 centimeters by 9 centimeters? Include correct units with your solution.

OpenStudy (anonymous):

\[5.04m ^{3}\]

OpenStudy (anonymous):

how did you get that answer

OpenStudy (anonymous):

That's 4 times 14 times 9 = 504cm^3=5.04m^3.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how did you get 5.04 from 504

OpenStudy (anonymous):

I changed the units from cm^3 to m^3. Since 1m = 100cm, 1m^3 = 100cm^3 so you can divide 504cm^3 by 100 to get 5.04m^3.

OpenStudy (anonymous):

ohh. ok.

OpenStudy (anonymous):

No - sorry. My mistake. There are 1,000,000cm^3 in 1m^3. The correct answer is just 504cm^3.

OpenStudy (anonymous):

100cm x 100cm x 100cm = (100)^3 x (cm)^3 = (100cm)^3 = 1m^3.